### v3

# FedSEA-LLaMA: A Secure, Efficient and Adaptive Federated Splitting Framework for Large Language Models

Abstract
Private data holds promise for improving LLMs due to its high quality, but its scattered distribution across data silos and the high computational demands of LLMs limit their deployment in federated environments. To address this, the transformer-based federated split models are proposed, which offload most model parameters to the server (or distributed clients) while retaining only a small portion on the client to ensure data privacy. Despite this design, they still face three challenges: 1) Peer-to-peer key encryption struggles to secure transmitted vectors effectively; 2) The auto-regressive nature of LLMs means that federated split learning can only train and infer sequentially, causing high communication overhead; 3) Fixed partition points lack adaptability to downstream tasks. In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated splitting framework based on LLaMA2. First, we inject Gaussian noise into forward-pass hidden states to enable secure end-to-end vector transmission. Second, we employ attention-mask compression and KV cache collaboration to reduce communication costs, accelerating training and inference. Third, we allow users to dynamically adjust the partition points for input/output blocks based on specific task requirements. Experiments on natural language understanding, summarization, and conversational QA tasks show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× speedups in training and inference. Further analysis of privacy attacks and different partition points also demonstrates the effectiveness of FedSEA-LLaMA in security and adaptability.
Codehttps://github.com/TAP-LLM/SplitFedLLM/tree/main/FedSEA-LLaMA
Extended versionhttps://arxiv.org/abs/2505.15683
Introduction
While the success of Large Language Models (LLMs) has largely stemmed from their ability to process vast amounts of public data (Kaplan et al. 2020), the true potential of these models lies in their application to high-quality private data. Private datasets often contain more granular and specific insights about individuals or organizations, typically collected through more controlled and precise methodologies (Yang et al. 2024). Additionally, private data tend to be more timely, targeted, and unique, thus providing models with semantic contexts and environmental settings that more closely mirror real-world scenarios (Ye et al. 2024). Therefore, effectively incorporating private data will further advance the performance of LLMs.
[…]
The distributed private data silos and the significant computational demands of LLMs pose substantial challenges for their deployment in federated environments. Private data is often isolated in local silos, such as mobile devices (Ghimire and Rawat 2022), enterprise servers (Huang et al. 2024), or medical institutions (Yuan et al. 2024), as centralized storage and processing of such data may raise privacy concerns and regulatory risks (Chen et al. 2024). Federated Learning (FL) (Stremmel and Singh 2021) offers a promising solution by enabling local training on user devices and only sending model parameters for aggregation. They often leverage Adapters (Houlsby et al. 2019) or LoRA (Hu et al. 2022) as transmitted parameters, or employ collaborative training between large and small models (Hinton, Vinyals, and Dean 2015) to fine-tune LLMs. However, in extremely resource-constrained scenarios (e.g., less than 3.5GB of GPU memory), fine-tuning the LLMs remains infeasible. Therefore, how to effectively deploy LLMs in FL frameworks while ensuring privacy and scalability in resource-constrained scenarios remains an open research problem.
To address this, transformer-based federated splitting models (Zheng et al. 2024; Su et al. 2024) have been proposed to adapt LLM fine-tuning to resource-constrained scenarios. Titanic (Su et al. 2024) handles limited server resources by partitioning LLM across clients and enabling collaborative training. FL-GLM (Zheng et al. 2024) offloads most model parameters to a resource-ample server while retaining the input and output layer on the client. As shown in Figure 1a, FL-GLM divides Transformers into three parts, i.e., the first input block and the last output block on the client, while the middle blocks on the server. The forward propagation (FP) path is client-to-server-to-client with peer-to-peer key encryption. After the client calculates loss, gradients backpropagate (BP) from client-to-server-to-client with the same encryption. Although FL-GLM provides a potentially viable solution for clients in resource-constrained scenarios, it still faces several limitations: 1) Privacy Leakage: during forward and backward propagation, activation values and gradients from the input blocks are transmitted, which can be exploited to reconstruct original input data through reverse engineering (Asnani et al. 2023). 2) Communication Cost: LLMs are auto-regressive, requiring sequential token generation based on prior tokens. In federated split learning, this necessitates a full forward and backward pass with client-server communication for each token, greatly increasing communication rounds and bandwidth usage. 3) Lack of Adaptability: Fixed partition points of FL-GLM lack flexibility to downstream tasks and hardware constraints, as different devices vary in their capacity to handle workloads and parameter sizes.
In this paper, we introduce FedSEA-LLaMA, a Secure, Efficient, and Adaptive Federated split framework based on LLaMA2, as shown in Figure 1b. To Enable Secure Propagation, we add Gaussian noise for the hidden states of input blocks and send these noised states to the server for further processing. Through both gradient analysis and experimental verification, this noise ensures privacy and security during subsequent training and inference stages. To Reduce Communication Cost, we introduce the attention-mask compression and collaborative KV cache mechanisms, minimizing overhead while preserving the integrity of auto-regressive inference. By synchronously maintaining cache states on both client and server, the originally cross-device transmission is transformed into dynamic attention-mask generation, reducing communication from megabytes (MB) to bytes (B). To Achieve Adaptive Partition, we design a dynamic partition strategy allowing users to flexibly adjust the number of input and output blocks according to specific downstream tasks and available computational resources. This adaptability enables an optimal balance between performance and efficiency.
Experiments on natural language understanding (NLU), summarization, and conversational QA show that FedSEA-LLaMA maintains performance comparable to centralized LLaMA2 and achieves up to 8× inference speedups. Through inference acceleration, the average token generation time of FedSEA-LLaMA has been reduced by 87.6%, and the speed gain will further increase with the growing length of context. Additional analysis of privacy attacks and various partition points further highlights FedSEA-LLaMA’s effectiveness in terms of security and adaptability.
The innovations in this paper are as follows:
  • We design a federated split framework tailored for resource-limited scenarios, featuring adaptive partitioning and client-side Gaussian noise injection to enable privacy-preserving training and inference of LLMs.
  • We leverage collaborative KV caching and dynamic attention-mask generation to drastically reduce communication overhead from megabytes to bytes during auto-regressive inference, achieving up to 8x speedups.
  • Experimental results on three different tasks show that our FedSEA-LLaMA achieves performance comparable to centralized LLaMA2, and validate the security, efficiency, and adaptability of our framework.
Related Works
[…]
Federated Learning for LLMs
Traditional Federated Learning (FL) frameworks face scalability challenges with large language models (LLMs) due to high communication and computational costs (McMahan et al. 2017; Stremmel and Singh 2021; Ji et al. 2019). Recent federated LLM frameworks (Fan et al. 2023; Kuang et al. 2024; Ye et al. 2024) address this by using instruction tuning, Adapters, and LoRA to reduce parameter updates, lowering communication overhead. However, transmission demands often still exceed system limits. To mitigate this, gradient compression (Wu et al. 2024b; Shu et al. 2024) and quantization (JianHao et al. 2024) techniques have been applied. Structural pruning based on Fisher Information (Liu et al. 2024) further narrows updates to key layers, reducing model size without performance loss. Knowledge distillation (Wu et al. 2024a; Fan et al. 2024, 2025) offers another solution, using smaller proxy models to transfer knowledge with minimal communication. Client heterogeneity remains a major challenge (Zhang et al. 2024), prompting methods like LoRA-based stacking (Bai et al. 2024; Wang et al. 2024) to align models from diverse clients and improve fairness.
Despite advances in efficient fine-tuning and parameter reduction for federated LLMs, resource-constrained clients still face challenges with independent fine-tuning, particularly under compression overhead. Split learning (Thapa et al. 2022) offers a promising alternative in such settings.
Split Learning for LLMs
Split learning is an emerging decentralized training approach designed for resource-constrained settings (Abedi and Khan 2024; Rahman et al. 2020; Matsubara and Levorato 2021), especially mobile or GPU-limited clients. It works by dividing a neural network into sub-networks, each processed on a different device for collaborative learning. To address the training of sequential data in LLMs, FedBERT (Tian et al. 2022) introduces an innovative federated learning framework designed to pre-train language models in a split architecture, addressing the challenge of limited computational capacity on client devices. Although this design enables more efficient distribution of training workloads, it comes with drawbacks, such as increased communication overhead, fixed partition points, and susceptibility to privacy threats like embedding gradient attacks.
Recently, Su et al. (2024) extends the scope of split learning by exploring scenarios with heterogeneous resources, particularly considering a case where the server lacks sufficient computational capacity. In their approach, the LLM is automatically divided into $k$ partitions, each assigned to one of $k$ clients, while the server merely performs aggregation. Although this strategy alleviates server-side computation, it considerably increases the overall training and inference time. Furthermore, repeated transmission of hidden states increases the risk of single points of failure and introduces potential security vulnerabilities. Zheng et al. (2024) propose FL-GLM to place the input and output blocks locally on client devices, while the remaining primary model parameters are centralized on a server with ample computational resources. And then, they employ key encryption during client-server communication to prevent reverse engineering attacks from peer clients. But the peer-to-peer key encryption is not enough because the peer clients and servers can collaborate to infer the clients’ private data.
Our work diverges from these studies by enhancing both training/inference efficiency, adaptive partition points, and data privacy for distributed LLMs through optimized algorithmic designs and communication strategies.
Model
In this section, we present a detailed description of the FedSEA-LLaMA framework, as shown in Figure 2, which comprises three components: privacy preservation, inference acceleration, and adaptive partitioning. Following FL-GLM, we adopt split learning to divide the LLaMA into three parts: the local client stores the first p{p} blocks($Blocks_{A}$) and the last q{q} blocks($Blocks_{C}$), while the remaining k{k} blocks($Blocks_{B}$), comprising the majority of the model parameters, are hosted on the server.
Privacy Preservation
In distributed training of LLMs, the split learning paradigm partitions the entire model into multiple sub-models deployed across different physical nodes. As a result, the hidden states and gradients must be transferred across nodes via the network. Unlike traditional federated learning, which only shares gradients, this approach exposes mid-layer activations and their gradients, greatly increasing the risk of adversaries reverse-engineering user data and challenging the data protection framework of federated split LLMs.
[…]
Threat Model Definition
We design an extreme attack scenario to evaluate the robustness of FedSEA-LLaMA. Inspired by Pasquini, Ateniese, and Bernaschi (2021), we construct a model inversion attack in a multi-client federated setting, as illustrated in Figure 3a, where a malicious client colludes with an honest-but-curious server. The malicious client shares its private data in plaintext with the server, allowing the server to train an attack model locally. The goal of the attack model is to reconstruct the original input text from the hidden states transmitted by the client. During this process, the server does not tamper with the hidden states sent to other honest clients, thereby performing a stealthy inversion attack against their private data without awareness.
Concretely, we consider an adversary that has full knowledge of the client-side model architecture and the number of layers deployed on the client, but does not have access to the actual model parameters. Thus, the honest-but-curious server locally initializes a decoder with the same architecture and depth as the client’s model and uses it as the attack model. Without loss of generality, we consider a two-client FedSEA-LLaMA setting, where the server receives hidden states from both $client_{0}$ and $client_{1}$. During each training step, the server first performs a normal forward pass through $Blocks_{B}$ to obtain the intermediate hidden states $h_{B}$. It then feeds the hidden states from the malicious $client_{0}$ into the attack model and uses the corresponding private data $D_{\text{priv}}$ as supervision. Afterward, $h_{B}$ is split and sent back to the respective clients, allowing training to proceed as usual without disrupting the federated process.
Defense with Gaussian Noise Injection
We explore two defense strategies to protect the transmission in FedSEA-LLaMA: adding noise to gradients and adding noise to forward-pass hidden states. The first strategy introduces random Gaussian noise into the initial gradient during backpropagation (i.e., the gradient $\nabla h_{B}$), while the second applies Gaussian noise to the client’s first forward-pass output (i.e., the hidden state $h_{A}$). Empirical results reveal that federated split LLMs are highly sensitive to gradient perturbations, so injecting noise into gradients often leads to instability in federated training, with the loss frequently diverging to NaN after several training steps. In contrast, perturbing the hidden state $h_{A}$ during the forward pass yields more stable training. Therefore, FedSEA-LLaMA adopts the latter to enhance robustness while preserving privacy.
Next, we derive how the noise introduced in the forward pass can also propagate to the gradients during backpropagation. Without loss of generality, we apply noise perturbation to $h_{1}$ in the theoretical proof, although in practice we apply noise perturbation to $h_{A}$. The proof will be similar in both cases. Figure 3b illustrates the schematic of noise injection during the forward propagation of LLMs. Let $W_{n}$ denote the weight matrix of the down-projection fully-connected layer within the final MLP layer, and $W_{n+1}$ denote the weight of the output linear layer. In the noise-free scenario, the hidden states $h_{1}$ undergo forward propagation:
$h_{2}$ $=\mathcal{A}(h_{1})+h_{1},$ (1)
$h_{3}$ $=\text{RMSNorm}(h_{2}),$ (2)
$h_{4}$ $=\mathcal{M}(h_{3})+h_{2},$ (3)
$=W_{n}[Silu(gate_{proj}(h_{3}))\cdot up_{proj}(h_{3})]+h_{2},$ (4)
$h_{5}$ $=W_{n+1}\cdot\text{RMSNorm}(h_{4}),$ (5)
$\hat{y}$ $=\text{softmax}(h_{5}),$ (6)
[…]
where $\text{RMSNorm}(*)$ is the layernorm function, $\mathcal{A}$ is the attention layer, $\mathcal{M}$ is the MLP layer.
During backpropagation, according to (Clark 2017), the gradient of $W_{n}$ is:
$\frac{\partial L}{\partial W_{n}}$ $=\frac{\partial L}{\partial h_{5}}\cdot\frac{\partial h_{5}}{\partial h_{4}}\cdot\frac{\partial h_{4}}{\partial W_{n}},$ (8)
$\frac{\partial L}{\partial h_{5}}$ $=(\hat{y}-y),$ (9)
$\frac{\partial h_{5}}{\partial h_{4}}$ $=W_{n+1}\cdot J_{\text{RMSNorm}}(h_{4}),$ (10)
$\frac{\partial h_{4}}{\partial W_{n}}$ $=[Silu(gate_{proj}(h_{3}))\cdot up_{proj}(h_{3})]^{T}.$ (11)
According to Eq. 1-11, the gradient of $W_{n}$ is closely related to the value of $h_{1}$. Thus, when Gaussian noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$, perturbations are consequently introduced into the gradients of parameters, thereby influencing subsequent gradient computations. In other words, injecting noise into the forward propagation process can protect the hidden states and their gradients transmitted during the training of the splitting LLMs.
Inference Acceleration
The inference time of LLMs is crucial for the efficient and timely execution of complex tasks. To enhance the efficiency and usability of distributed LLMs, our proposed FedSEA-LLaMA framework incorporates attention-mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
Attention-Mask Compression
FedSEA-LLaMA transmits only the minimal information necessary for effective model training and inference. In particular, it reduces redundancy in attention-mask transmissions compared to FL-GLM. Since the generation of attention-mask remains consistent across transformer blocks within each training iteration and depends solely on dataset padding, transmitting full attention-mask tensors $A_{mask}\in[0,-\infty]^{batch\_size\times seq\_len\times seq\_len}$ results in unnecessary communication overhead during training and inference. Instead, FedSEA-LLaMA transmits only essential metadata (seq_lenseq\_len, pad_lenpad\_len), indicating padding locations, thereby significantly decreasing bandwidth usage.
Collaborative KV Cache Mechanism
In traditional distributed inference of LLMs, each inference step involves transmitting all previously generated tokens from the client to the server. This design leads to two major inefficiencies: (1) the computational overhead increases linearly with the growing input length, since the attention mechanism must recompute key and value (KV) representations for all past tokens; and (2) the repeated transmission of the entire context between client and server introduces substantial communication latency, especially for long sequences.
To address these issues, we propose a collaborative KV cache mechanism tailored for distributed inference. Instead of recomputing and retransmitting all past tokens at each step, our approach caches the attention KV states locally on both the client and server sides. Since these KV states remain static across the autoregressive generation process, subsequent tokens can reuse the cached representations. Built on this mechanism, FedSEA-LLaMA enables highly efficient distributed inference by significantly reducing both computational and communication overhead. Experimental results show that this strategy leads to substantial improvements in inference speed in scenarios with long user queries.
Adaptive Partition
To flexibly accommodate specific task requirements and hardware constraints, we propose an adaptive splitting strategy that enables users to adjust the number of model input and output blocks, thereby achieving an optimal balance between performance and efficiency. Based on this strategy, we explore the performance of downstream tasks when the client holds the first blocks (with $Blocks_{A}$=1,2,3) and the last blocks (with $Blocks_{C}$=1,2,3) of the model. The results indicate that the depth of the model split does not significantly impact performance on downstream tasks, while affording a more flexible allocation of computational loads between the server and the clients. Furthermore, it corroborates the efficiency of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
Experiments
We evaluate FedSEA-LLaMA on NLU, summarization, and conversational QA to assess its performance against centralized training while tackling latency and privacy via decentralized learning. More experiments, including parallel training and training prompts, can be found in the Appendix.
Model ReCoRD COPA WSC RTE BoolQ WiC CB MultiRC CoQA Xsum F1/Acc. Acc. Acc. Acc. Acc. Acc. F1 F1a/EM F1/EM rouge-1/rouge-2 $\text{T5}_{\text{large}}$ (Raffel et al. 2020) 85.7/85.0 78.0 84.6 84.8 84.3 71.6 96.4 80.9/46.6 - 40.9/17.3 $\text{BART}_{\text{Large}}$ (Lewis et al. 2020) 88.3/87.8 60.0 65.4 84.5 84.3 69.0 90.5 81.8/48.0 - 45.1/22.3 $\text{GLM}_{\text{RoBERTa}}$ (Du et al. 2022) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (Zeng et al. 2022) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 FL-GLM (Zheng et al. 2024) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 62.7/49.0 37.0/11.9 LLaMA2-7B (Touvron et al. 2023) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
Table 1: Performance Comparison on SupleGLUE, CoQA and Xsum datasets between FedSEA-LLaMA and baselines.
Experimental Settings
In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines, and parameter settings for FedSEA-LLaMA.
Dataset
To evaluate model generalization across diverse NLP tasks, our experiments incorporate three distinct tasks:
  • The SuperGLUE benchmark (Wang et al. 2020) evaluates eight distinct NLU tasks, from common sense reasoning to fine-grained semantic interpretation.
  • The CoQA dataset (Reddy, Chen, and Manning 2019) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues across seven domains. Answers are free-form text, with nearly half of the questions requiring coreference resolution and pragmatic reasoning, simulating real-world dialogue coherence.
  • The XSum dataset (Narayan, Cohen, and Lapata 2018) evaluates abstractive summarization via extreme compression of BBC news articles into single-sentence summaries. It tests the model’s ability to abstract and condense salient information across diverse topics.
Metrics
We evaluate model performance using standard task-specific metrics. For NLU tasks such as COPA, WSC, etc, we report Accuracy, which measures the proportion of instances where the model’s predicted label exactly matches the ground-truth label.
For conversational QA tasks, we use Exact Match (EM) and F1 score to evaluate model performance (Yatskar 2019; You et al. 2022). EM measures the percentage of predictions that exactly match the ground truth, accounting for punctuation and case, while F1 reflects the harmonic mean of precision and recall, rewarding partial overlaps.
For summarization tasks, we use standard ROUGE metrics (Lin 2004; Liu et al. 2021; Fang et al. 2022), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
Baselines
We evaluate FedSEA-LLaMA based on the LLaMA2-7B model (Touvron et al. 2023), an open-source pre-trained model developed by Meta and FL-GLM. We additionally include several strong baselines: $\text{T5}_{\text{large}}$ (Raffel et al. 2020), $\text{BART}_{\text{Large}}$ (Lewis et al. 2020), $\text{GLM}_{\text{RoBERTa}}$ (Du et al. 2022) and ChatGLM-6B (GLM et al. 2024).
Parameter Settings
The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40. The implementation is based on the Flower (Beutel et al. 2020) framework, which orchestrates the transfer of intermediate data between clients and the server. Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset. All other experiments are conducted with FP16 precision and a batch size of 2. The random seed is consistently set to 42. Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters. For all LoRA fine-tuning experiments in this paper, the target modules are set to q_proj, v_proj, k_proj, and o_proj.
Experimental Results
FedSEA-LLaMA Performance
We compared the performance of FedSEA-LLaMA with centralized LLaMA2-7B on various tasks, including NLU, summarization generation, and conversational QA, as shown in Table 1. Experimental results show that FedSEA-LLaMA achieves comparable performance to LLaMA2-7B, with metrics approaching or matching across multiple tasks, demonstrating its effectiveness. Meanwhile, compared to other baselines, FedSEA-LLaMA achieves better performance on datasets such as Record, MultiRC, and Xsum, but performs poorly on the RTE dataset. This is speculated to be related to the distribution difference between RTE and the data used for LLaMA2 pre-training, as well as suboptimal hyperparameters. The comparison of their GPU memory usage is shown in Figure 4. The result shows that FedSEA-LLaMA significantly reduces GPU memory consumption on the client side. Both models were fine-tuned on the Record dataset with a batch size of 2, using float16 precision and a LoRA rank of 8. While centralized LLaMA2-7B occupies 28.2G of GPU memory, FedSEA-LLaMA client requires only 3.4G, representing a reduction of up to 87.9% in memory requirements.
Figure 4: Computation load on client side of centralized LLaMA2-7B and FedSEA-LLaMA on ReCoRD dataset.
(a) Loss Curves for Noise Scales
($Blocks_{A}$=1, $Blocks_{C}$=1)
(b) Inference Speed Comparison
(w/ vs. w/o Attention-Mask).
(c) Inference Speed Comparison
(w/ vs. w/o KV Cache).
Figure 5: (a)Different noise injection on MultiRC dataset. (b)Ablation study on attention-mask. (c)Ablation study on KV cache.
Effects of Noise Scales
Since noise injection into the hidden states during forward propagation inevitably impacts the performance of distributed LLMs, we conducted extensive experiments to investigate how noise levels affect model performance. The corresponding experimental results are presented in Figure 5(a) and Table 2. Figure 5(a) demonstrates that when the numbers of blocks in $Blocks_{A}$ and $Blocks_{C}$ are consistent, the convergence speed of FedSEA-LLaMA decreases as the magnitude of injected noise increases. When the average absolute value of the noise is small (e.g., below 0.02), the loss curve remains close to that of the noise-free baseline. Table 2 indicates that under fixed noise levels, increasing the number of blocks in $Blocks_{A}$ consistently improves downstream performance in terms of both F1 and EM scores. Notably, when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness.
Noise Scale $Blocks_{A}$ = 1 $Blocks_{A}$ = 2 $Blocks_{A}$ = 3 0.0 77.48/40.82 77.82/39.24 77.89/41.34 0.02 77.89/38.09 77.87/39.66 78.61/42.39 0.05 74.57/29.17 77.07/38.41 78.95/41.03 0.1 59.11/8.18 72.45/23.71 75.68/34.10
Table 2: Different noise scales and blocks number of $Blocks_{A}$ analysis(F1a / EM) on MultiRC dataset.
Model Reverse Attack
To assess how split points and noise scale affect the security of FedSEA-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth ($Blocks_{A}$ = 1, 2, 3) and noise (scale = 0, 0.02, 0.5). Attacks use full-parameter fine-tuning (lr=2e-5), with reconstruction quality measured by BLEU-4 and ROUGE-2. FedBert reports a Rouge-2 of 7.053 and a Bleu-4 of 28.57. Experiment results of FedSEA-LLaMA are shown in Table  3. From the results, we can see that when the client holds only the embedding layer (similar to FedBERT), the attack model achieves relatively high BLEU and ROUGE scores. However, under the FedSEA-LLaMA framework, where the client also retains a portion of the LLM blocks in addition to the embedding layer, all attack metrics approach zero. These findings demonstrate that FedSEA-LLaMA effectively mitigates model inversion attacks, thereby preserving the privacy of client-side data.
Noise scale $Blocks_{A}$=1 $Blocks_{A}$=2 $Blocks_{A}$=3 0.0 0.051/0.003 0.056/0.003 0.069/0.016 0.02 0.050/0.002 0.052/0.003 0.084/0.044 0.05 0.050/0.003 0.051/0.002 0.067/0.004
Table 3: Security analysis (Rouge-2 F1 / BlEU-4) of FedSEA-LLaMA on Record dataset.
Efficiency Optimization
Figure 5(b) shows a comparison of the time consumption between transferring the attention-mask and transferring only the actual input length during distributed training. The bar chart represents the average time spent per token in FedSEA-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second(context length=500). We can see that transferring the compressed information of the attention-mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
As shown in Figure 5(c), FedSEA-LLaMA demonstrates superior performance compared to FL-GLM in distributed inference, achieving over 7× speedup. Without the KV cache, the inference speed decreases significantly with longer queries, due to the increasing number of K and V vectors. In contrast, with the KV cache technique, the client only needs to send the intermediate vector of the current single token to the server, while reusing previously generated K and V vectors. This effectively avoids redundant computations. When the query length reaches 4000, the inference speed of FedSEA-LLaMA with KV cache is 8.2 times that without it. As query length increases further, this advantage continues to expand, greatly alleviating the computational burden and operational latency of distributed LLMs.
Adaptive Splitting
The choice of model partitioning has a significant impact on the computational burden at the client side. To evaluate whether FedSEA-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different numbers of blocks in $Blocks_{A}$ and $Blocks_{C}$ ($Blocks_{A}$ = 1, 2, 3 and $Blocks_{C}$ = 1, 2, 3). This setting reflects realistic constraints where clients possess limited computational resources. Table 4 shows F1a/EM scores across different partitioning schemes. Performance variations across partitions are minimal, indicating that FedSEA-LLaMA is robust to different split configurations. This flexibility allows resource-constrained clients to adjust block allocation based on their downstream task requirements and resource availability, without significant accuracy loss.
$Blocks_{A}$ $Blocks_{C}$ = 1 $Blocks_{C}$ = 2 $Blocks_{C}$ = 3 1 80.34/44.60 80.63/46.59 80.90/47.11 2 80.41/45.86 80.62/45.96 79.94/45.02 3 81.17/47.22 80.74/46.59 80.20/45.33
Table 4: Different model partition analysis(F1a/EM) on MultiRC dataset. No Noise is added to the hidden states.
Conclusion
In this work, we propose FedSEA-LLaMA, a novel federated split learning framework tailored for LLMs. By securely offloading the majority of model computation to a central server while maintaining privacy-sensitive components on local clients, FedSEA-LLaMA addresses the core limitations of traditional federated split learning approaches. Through secure end-to-end propagation with noise injection, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FedSEA-LLaMA demonstrates significant advances in security, efficiency, and adaptability. This federated split LLMs framework not only keeps users’ private data confined to their local devices but also significantly shortens training and inference time, making it better suited for handling the scale and complexity of LLMs. In future work, we aim to support more base models and deploy the FedSEA-LLaMA framework in scenarios where data privacy is a critical concern.
Acknowledgements
This work was funded by the National Natural Science Foundation of China (NSFC) under Grants No. 62406013,No.62388101, the Beijing Advanced Innovation Center Funds for Future Blockchain and Privacy Computing(GJJ-24-034), and the Fundamental Research Funds for the Central Universities.
Appendix A Noise Added to Hidden States
(a) Loss Curves for Noise Scales
($Blocks_{A}$=1, $Blocks_{C}$=1)
(b) Loss Curves for Noise Scales
($Blocks_{A}$=2, $Blocks_{C}$=1)
(c) Loss Curves for Noise Scales
($Blocks_{A}$=3, $Blocks_{C}$=1)
Figure 6: FedSEA-LLaMA: Noise Injection across different number of blocks in $Blocks_{A}$ on MultiRC dataset.
The magnitude of noise injection significantly influences the stability and convergence speed of model training. Given that noise is directly injected into the hidden state $h_{A}$, the relative magnitude between the noise and $h_{A}$ naturally affects the model’s performance. To quantify this effect, we first fine-tune the FedSEA-LLaMA model on the MultiRC (Khashabi et al. 2018) with a batch size of 2 and a learning rate of $1\times 10^{-5}$, analyzing the average absolute magnitude of all elements within the hidden state $h_{A}$. According to Figure 7, the average magnitude of the hidden state $h_{A}$ is approximately 0.02 ($num_{Blocks_{A}}$=1). The average magnitude of the hidden state $h_{A}$ increases with the number of blocks in $Blocks_{A}$.
Based on this observation, we subsequently inject noise at various scales around this value during training to investigate how the noise magnitude affects both the convergence speed and downstream task performance. Training is conducted for 5,000 steps with FP16 precision. We use LoRA with a rank of 8 and a scaling factor $\alpha=16$. The learning rate is 1e-5, and the batch size is 2. Corresponding experimental results are presented in Figure 6(a)-6(c) and Table 2. To assess the sensitivity to the noise scale when $Blocks_{A}$=3, we evaluate performance across a range of noise scales (from 0.1 to 0.4). As shown in Table 5, model performance decreases as the noise scale increases. This suggests that while moderate noise injection is tolerable, excessive noise harms representation quality.
Noise Scale F1a EM
0.1 75.68 34.10
0.15 71.21 26.82
0.2 59.08 16.47
0.3 55.79 7.45
0.4 57.68 6.09
Table 5: Performance under different noise scales when $Blocks_{A}$=3
A horizontal comparison reveals that, under the same noise level, increasing the number of blocks in $Blocks_{A}$ consistently accelerates convergence. Table 2 indicates that when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness. This can be explained by the average magnitude shown in Figure 7: as the depth of $Blocks_{A}$ increases, the ratio between the average absolute value of the hidden state $h_{A}$ and the noise amplitude also increases, thereby improving the signal-to-noise ratio (SNR). This enhanced SNR is precisely what contributes to the better convergence and stronger robustness observed with a larger number of $Blocks_{A}$.
Figure 7: Avg magnitude of hidden state $h_{A}$ ($Blocks_{C}$ =1) on MultiRC dataset.
Appendix B Prompt
Table 6 illustrates examples of cloze-style questions and corresponding answers derived from each task. (1)ReCoRD tests reading comprehension with common sense reasoning by resolving ambiguous entities in passages. (2)COPA assesses causal reasoning by choosing the more likely cause or effect for a given premise. (3)WSC focuses on pronoun disambiguation requiring deep contextual understanding. (4)RTE evaluates textual entailment by determining if one sentence follows, contradicts, or is unrelated to another. (5)BoolQ involves yes/no questions answered using evidence from context passages. (6)WiC tests word sense disambiguation by checking if a word has the same meaning in two sentences. (7)CB deals with entailment in complex sentences with nuanced reasoning. (8)MultiRC features multi-answer questions based on multi-sentence contexts, requiring integration of dispersed information.
Table 7 illustrates examples of fine-tuning templates of CoQA and XSum.
Dataset Task Cloze Question Answers ReCoRD Question answering [passage p] [cloze question q] Answer candidates COPA Causal reasoning “[choice c1]” or“[choice c2]”? [premise p], so [M]. c1/c2 WSC Coreference resolution [sentence s] The pronoun ‘*p*’ refers to [M]. Noun n RTE Textual entailment “[hypothesis h]”? [M] “[premise p]” “yes”/“no” BoolQ Question answering [passage p]. Question: q? Answer:[M]. “yes” / “no” WiC Word sense disambiguation “[sentence s1]”/“[sentence s2]”Similar sense of [word w]? [M]. “yes”/“no” CB Textual entailment “[hypothesis h]”? [M], “[premise p]” “yes”/“no”/“maybe” MultiRC Question answering [passage p]. Question: q? Is it [answer a]? [M]. “yes”/“no”
Table 6: Cloze questions and answers for the 8 SuperGLUE tasks
Dataset Task Fine-tuning template
CoQA conversational QA
<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
You need to answer my question base on the provided story.<</SYS>><</SYS>>
The story is {story}.My question is {question}.[/INST]{answer}</s></s>
XSum summarization generation
<s><s>[INST] <<SYS>><<SYS>>You are a helpful assistant.
You need to summarize an abstract base on the provided passage.<</SYS>><</SYS>>
The passage is {passage}.[/INST]{summarization} </s></s>
Table 7: Fine-tuning templates for CoQA and XSum
Appendix C Parallel Training
[…]
Table 8: Comparison of training time between different training strategies on ReCoRD dataset.
(a) Client-batch parallel training.
(b) Server-hierachical parallel training.
Figure 8: Multi-client Parallel Training of FedSEA-LLaMA. In (a), the server receives forward-pass states from multiple clients simultaneously, concatenates them into a batch, and performs joint forward propagation. Then, the server splits hidden state $h_{B}$ and sends them back to different clients. In (b), each $client_{i}$ is paired with a corresponding $sub\_server_{i}$. After several training steps, the central server aggregates all sub-servers’ models using FedAvg and distributes the updated global server model back to each $sub\_server_{i}$.
In multi-client settings, the FedSEA-LLaMA framework accommodates three training paradigms: sequential training, client-batch parallelism, and server-hierarchical parallelism, each designed for different system configurations and resource availabilities.
Sequential Training
The client begins by passing hidden states through its input blocks $Blocks_{A}$ and forwards the result to the server. The server processes the input through $Blocks_{B}$ and returns the output states. The client then computes the loss with $Blocks_{C}$, sends the gradients to the server for model updates, and receives the updated gradients to adjust its input blocks. This process continues until all clients have completed their respective training phases for the current round. Although this method is inherently slower due to its serialized nature, it significantly reduces the computational burden on the server, making it a suitable choice for scenarios where server-side resources are limited.
Client-batch Parallelism
The client-batch parallel approach enhances training throughput by expanding the effective batch size at the server, as shown in Figure  8(a). During each training round, the server collects intermediate hidden states from all participating clients. If each of the $M$ clients processes a local batch of size $b$, the hidden states sent to the server will each have the shape $(b,\textit{seq\_length},\textit{hidden\_size})$. The server concatenates these hidden states along the batch dimension, forming a combined tensor of shape $(b\times M,\textit{seq\_length},\textit{hidden\_size})$. This aggregated batch is then used to perform a joint forward and backward pass on the server model, allowing it to compute gradients more efficiently and update shared parameters.
Server-hierarchical Parallelism
In a server-hierarchical parallel setup, the central server spawns multiple sub-server model instances, each dedicated to a specific client, illustrated in Figure  8(b). These instances run in parallel, enabling simultaneous local training for all clients. Typically, after a fixed number of local updates, the central server synchronizes the training progress by aggregating model parameters across all sub-servers. This is usually done via weighted averaging of both server-side and client-side model parameters. This strategy enables significant improvements in training efficiency and scalability, especially in environments where the server has access to multiple GPUs or multi-threaded execution capabilities.
For communication cost analysis, we randomly select 1,000 samples from the ReCoRD dataset. All training uses a batch size of 1 with FP16 precision for efficiency. LoRA is applied with a rank of 8 and a scaling factor of 16, targeting attention modules during fine-tuning. The model is partitioned with $Blocks_{A}$=1 and $Blocks_{C}$=1. Each experiment is repeated five times, and we report the mean and standard deviation of the total communication time in Table 8. Results show that sequential training is the slowest, client-batch training is faster, and server hierarchical training is the fastest. Notably, when the server can handle larger batches, increasing the number of clients in client-batch training effectively reduces training time.
Appendix D Efficiency Comparison between FedSEA-LLaMA and FL-GLM
For the measurements of inference speed, inference memory usage, and per-token communication volume, we used the following prompt template:
{B_INST}{B_SYS}System: You are a helpful assistant. Always answer as helpfully as possible.{E_SYS}User: what can you do?{E_INST}
The total prompt length was 44 tokens. All inference experiments were conducted in FP16 precision with a maximum output length of 128 tokens. During inference, the client held one layer of model A and one layer of model C.
For the training memory usage experiments, we used the MultiRC dataset. Both model A and model C contained one layer. Unless otherwise specified, the default configuration was as follows: learning rate $1\times 10^{-5}$, LoRA rank =32=32, LoRA scaling factor =32=32, random seed =64=64, FP16 precision, and batch size =2=2. It is important to note that our training implementation actively clears unused variables in real time to minimize GPU memory usage. Therefore, the reported memory values correspond to the maximum observed usage within a one-minute monitoring window. All experiments were repeated three times with different random seeds (42, 64, and 128) to ensure the robustness and reproducibility of the results. The comparison results are shown in the Table 9.
Framework
Training Memory Usage
Inference Memory Usage
per-token communication
volume (Bytes)
Inference tokens
per second
FL-GLM 9810±0.00 6886±0.00 context length N × 10,335 0.75±0.03
FedSEA-LLaMA 5078±0.00 3674.00±0.00 8494 0.34±0.01
Table 9: Efficiency Comparison between FedSEA-LLaMA and FL-GLM
Appendix E Parameter Setting
FedSEA-LLaMA and Baselines
Different hyperparameters are explored to obtain the best performance of FedSEA-LLaMA and Centralized LLaMA. Specifically, we tune the LoRA rank (r) and scaling factor (alpha) as pairs, trying combinations including (4, 32), (8, 32), (16, 32), (32, 32), and (8, 16). For the learning rate (lr), values of 1e-5, 2e-5, and 3e-5 are tested. The final parameter settings were selected based on the model’s validation performance. The final parameters are shown in Table 10.
CoQA XSum MultiRC Record
FedSEA-LLaMA Lora r & alpha $(4,32)$ $(16,32)$ $(32,32)$ $(32,32)$
lr 1e-4 3e-5 1e-5 2e-5
Centralized-LLaMA Lora r & alpha $(8,16)$ $(32,32)$ $(32,32)$ $(8,32)$
lr 5e-5 3e-5 1e-5 3e-5
Table 10: Hyperparameters for FedSEA-LLaMA and Baselines
Different Noise Scales and Layers of $Block_{A}$ Experiment
We use a LoRA rank of 8 with a scaling factor of 16, a learning rate of 1e-5, and FP16 precision to investigate the effects of different noise scales and the number of blocks in $Blocks_{A}$ on downstream performance.
Ablation Study on Attention-Mask
The context length is set to 500, and the precision is set to FP16.
Ablation Study on KV Cache
The precision is set to FP16. Different context lengths are set to investigate the inference speed of FedSEA-LLaMA and FL-GLM.
Different Model Partition
All experimental results were obtained after 8,000 training steps with LoRA rank set to 8, scaling factor set to 16, a batch size of 1, FP16 precision, and a learning rate of 1e-5.